# FIITED: Fine-grained embedding dimension optimization during training for recommender systems

### v1

@@ 4.2 Dimension Optimization Algorithm
The dimension optimization procedure is detailed in Algorithm 1. Chunk utility values are updated in every training iteration. The utility values can simply record the running average of row access frequency, or it can be designed to incorporate more complex information, e.g., the L2 norm of gradients. Our preliminary results show that incorporating gradient information can lead to better model accuracy. Therefore, chunk utility in the $i$-th iteration is computed by $u(i)=\gamma u(i-1)+a(i)g(i)$, where $a(i)$ is the number of accesses to the chunk in iteration $i$, $g(i)$ is the L2 norm of the gradients computed for the chunk in iteration $i$, and $\gamma\in(0,1)$ is a decay parameter that reduces the influence of history utility values in order to capture dynamic changes in data characteristics.
@@
Embedding pruning occurs every $T$ iterations. Empirically, $T$ is set to around 1 hour worth of training data. During pruning, pruning thresholds are computed by first sorting the current utility values and then selecting the value at the desired pruning ratio. Instead of sorting all the utility values which can take a long time, a small number of utility values are randomly sampled. After obtaining the thresholds, embedding chunks with utility below the thresholds are pruned while the rest are retained in the model.
@@
4.2.1 Manual VS. adaptive per-chunk pruning ratios
Individual pruning ratios for each chunk can be selected either manually or adaptively according to data charcteristics. For manual selection, a pruning ratio $p_{k}$, $k=0,1,...,K-1$, for each chunk is selected empirically before training. For adaptive selection, a global pruning ratio $p$ is decided by the user and indicates the average pruning ratio across all the chunks. Utility values in all chunks are sorted together to decide one global pruning threshold. Individual pruning ratios for different chunks may be bigger or smaller than $p$, and may change over time, depending on chunk utility values computed at runtime. We evaluate both approaches in the experiments (Section 5).
4.2.2 Flexibility of the method
How to decide the utility value can affect the pruning behavior, which gives FIITED flexibility and expressiveness. For example, feature-wise statistics can be incorporated into row utility to make embeddings of the same sparse feature have similar dimensions. An extreme case is to only use feature-wise statistics for the utility computation, which will result in uniform-sized embeddings for the same sparse feature. In this paper, we choose to include gradient information in the row utility computation, which provides extra signals in deciding the importance of an embedding. Moreover, with such complex utility computation, chunks in the same embedding have different utility values, and as a result, per chunk pruning ratios can be automatically derived during training instead of manually chosen. On the other hand, setting the same utility for chunks in the same embedding (e.g., computing utility based solely on row access frequency) has unique benefits: combined with manually designed pruning ratios, such utility setting makes sure chunks are pruned in the same order in every embedding. For example, for K=2, if the pruning ratios are set to 50% and 90% for the two chunks and the chunks in the same embedding share the same utility, then the second chunk will always get pruned first in each embedding. The advantage of a fixed pruning order is to make sure the unpruned chunks stay in concentrated locations and that dot product between two embeddings lose less information. For example, for two embeddings with 1 chunk left, if the same pruning order is enforced, the remaining chunk should be in the same location and dot product between the two embeddings can capture the information in the unpruned chunks. However, if the pruning order is not the same, the remaining chunks may be in different positions and the dot product will simply return 0 without capturing any information. Selecting a suitable utility criterion as well as effective per-chunk pruning ratios is important in order to achieve satisfactory memory savings while maintaining the model quality.
4.3 System design
Although fine-grained in-training dimension pruning has multiple benefits, it faces one crucial practical challenge that is to realize memory saving during training: a straightforward implementation will result in many small fragments of free memory. To this end, we propose a new Virtually Hashed Physically Indexed (VHPI) embedding table design adapted from AdaEmbed (Lai et al., 2023), as illustrated in Figure 5. The original system design in AdaEmbed becomes a special case where the number of chunks is 1.
System Components The design mainly consists of a hash table, an embedding table and a chunk address manager. (1) VHPI hash table. Each hash table entry contains addresses of K embedding chunks, K utility values and a bit mask that indicates whether the chunks have been pruned. Since K is usually chosen to be small, the size of an entry is also small, which allows the hash table to have a large number of entries and a low collision rate. (2) VHPI embedding table. The embedding table stores embedding vector chunks coming from all sparse features. The size of the embedding table is set to a desired amount decided by the user, i.e., the pruning rate multiplied by the size of embeddings in an unpruned model. (3) Chunk address manager. It manages free addresses in the embedding table and maintains a free address stack which is updated with newly available chunk locations whenever a chunk is evicted during pruning.
Operations (1) Embedding access. To fetch an embedding, the sparse feature ID and the feature value are together hashed to obtain the index of an entry in the VHPI hash table. Unpruned chunks are identified by checking the K-bit mask, and their addresses are obtained from the entry. Embedding chunks are then fetched from the embedding table according to the addresses. (2) Embedding eviction. When an embedding chunk needs to be evicted, its address is passed to the chunk address manager, who pushes the address into the free address stack, and the K-bit mask in the hash table is updated accordingly. (3) Embedding allocation. To allocate an embedding chunk, the chunk address manager fetches the next available chunk location, which is then stored in the hash table entry. Embeddings at the allocated location are initialized and the K-bit mask in the hash table is updated.
Memory overhead The memory cost of the hash table depends on the value of K, with each entry requiring K location values, K utility-related values and K Boolean values. In practice, K is set to small values like 2, 4 or 8, so the size of the hash table is a lot smaller compared to the actual embedding table.
Latency overhead This approach takes (K+1) memory look-ups to fetch an entirely unpruned embedding, but the K look-ups in the embedding table can be carried out in independent threads and executed in parallel. Moreover, during distributed training, if the embedding table is sharded along the column dimension, then the number of chunks on each machine will be reduced to K/num_shards, further reducing the effect of the extra memory look-ups. In the best case, num_shards=K, and we will only have 1 extra memory lookup in each machine, which is the same as the row pruning method in AdaEmbed (Lai et al., 2023). The performance overhead of FIITED is evaluated quantitatively in Section 5.2.2.
@@ 4.3 System design
5 Experimental Results
5.1 Experimental setup
Baselines We compare FIITED with the following baselines: (1) AdaEmbed (Lai et al., 2023), a state-of-the-art in-training embedding pruning method that prunes entire embedding rows during training, as described in Section 4.1; (2) ESAPN (Liu et al., 2020), a state-of-the-art in-training EDS method that selects embedding dimensions by training additional policy networks that decide to enlarge or shorten embedding dimensions during training; (3) Mixed Dimension (MD) embeddings (Ginart et al., 2021), a state-of-the-art pre-training EDS method that selects per-feature embedding dimensions prior to training based on access frequency; (4) last but not least, the original unpruned DLRM model with uniform embedding dimensions, which will be denoted simply as DLRM in the comparison.
Models and datasets We evaluate FIITED on public models and datasets as well as private industry ones. A summary of the models and datasets can be found in Table 1. (1) Public models and datasets: We adopt an open-source DLRM framework 111 https://github.com/facebookresearch/dlrm to train a DLRM (Naumov et al., 2019) on Criteo Kaggle Display Advertising Challenge Dataset 222 https://ailab.criteo.com/ressources, Avazu Click-Through Rate Prediction Dataset 333 https://www.kaggle.com/c/avazu-ctr-prediction and Criteo Terabyte Dataset  [^4] https://labs.criteo.com/2013/12/download-terabyte-click-logs. We also modified this framework to train another open-source DLRM 555 https://github.com/zgahhblhc/ESAPN used by ESAPN (Liu et al., 2020) on the MovieLens-20M dataset 666 https://grouplens.org/datasets/movielens/20m. On the MovieLens-20M dataset, following ESAPN, we converted the multi-classification problem to a binary classification problem by viewing 4-star and 5-star reviews as positive, others as negative. (2) Industry models and datasets: Two production models are used in the evaluation. For ease of experimentation, both models are shrunk to 1/4 size by reducing the hash size of all sparse features by 75%. The reduced models, named Model A and Model B, have around 180GB and 430GB sparse feature embeddings, respectively, and both contain hundreds of sparse features. They are evaluated on data generated by real-world applications. Ten consecutive days of data are used as the training set, and the following one day’s data are used as the evaluation set.
Table 1: Models and datasets used for evaluation.
Model Total embedding size Dataset
DLRM 4.02GB Criteo-Kaggle
DLRM 1.13GB Avazu
DLRM 0.03GB MovieLens-20M
Model A (private) 180GB private
Model B (private) 430GB private
Implementation For evaluation on the public models and datasets, we implement FIITED as well as AdaEmbed (Lai et al., 2023) on top of an open-source DLRM framework (Naumov et al., 2019) based on PyTorch. Utility update is executed in parallel with training to reduce runtime overhead. The implementation will be open-sourced upon acceptance of the manuscript. For evaluation on the industry models and datasets, a proof-of-concept prototype design of FIITED is implemented on the company’s internal deep recommendation system code base. The prototype stores both pruned and unpruned embedding chunks, and executes pruning by setting the corresponding memory regions to zero. Gradients computed for pruned embedding chunks are also set to zero. For each row, a utility value is maintained as well as a bit mask indicating whether the chunks are pruned. If a chunk is re-allocated, the change will be reflected in the bit mask, and the subsequent gradient computation and weight updates will resume as usual. The correctness of the prototype is verified by checking the pruning decisions against the utility values as well as ensuring that the pruned chunks indeed remain zero. Although the prototype can only be used to evaluate the relationship between model quality and training memory usage and cannot reflect performance overhead, the performance overhead of FIITED can be measured on the public models and datasets (Section 5.2.2).
Training specifications For evaluation on the public models and datasets, the experiments are run on 1 GPU, and we compare two methods to select per-chunk pruning ratios: manually and dynamically. To manually specify the pruning ratio for each embedding chunk, a linear function is fitted to satisfy the desired average pruning ratio. We used a default sparse embedding dimension of 32 on Criteo Kaggle and Avazu datasets, 64 on Criteo Terabyte and MovieLens-20M datasets. The number of chunks K is set to 2 on all datasets for FIITED. For evaluation on the industry models and datasets, the experiments are run on 4 compute nodes with 8 GPUs per node. The average embedding pruning ratio varies from 30% to 85%. Per chunk pruning ratios are manually chosen and shown in Figure 6. The ratios are designed to roughly resemble a power law distribution, which was assumed by previous EDS work (Ginart et al., 2021). K varies between 4 and 8. The design also fits the observation that only a small number of “hot” embeddings require a long dimension Qu et al. (2022). For Model A, we also experimented with 2 different pruning ratio settings that have the same average ratio (75%), one is “steeper” with the ratios in different chunks span a larger range and the other is “flatter”, to investigate the effect of such a difference. Chunk utility is simply computed as a running average of access frequency in the experiments on industry models.
@@ 5.1 Experimental setup
Figure 6: Per chunk pruning ratio setup for industry models. Figure legends show the averaged embedding pruning ratio. The number of chunks varies from 4 to 8.
@@
Figure 7: Validation accuracy under different pruning ratios (62.5%, 75%, 87.5%, 93.75%, 99%) on the Criteo Kaggle dataset. Mixed Dimension by design cannot reach 99% pruning rate.
Figure 7: Validation accuracy under different pruning ratios (62.5%, 75%, 87.5%, 93.75%, 99%) on the Criteo Kaggle dataset. Mixed Dimension by design cannot reach 99% pruning rate.

### v2

@@ III-B Dimension Optimization Algorithm
The dimension optimization procedure is detailed in Algorithm 1. Chunk utility values are updated in every training iteration. The utility values can simply record the running average of row access frequency, or they can be designed to incorporate more complex information, e.g., the L2 norm of gradients. Our preliminary results show that incorporating gradient information can lead to better model accuracy. Therefore, chunk utility in the $i$-th iteration is computed by
$u(i)=\gamma u(i-1)+a(i)g(i)$ (1)
where $a(i)$ is the number of accesses to the chunk in iteration $i$, $g(i)$ is the L2 norm of the gradients computed for the chunk in iteration $i$, and $\gamma\in(0,1)$ is a decay parameter that reduces the influence of history utility values in order to capture dynamic changes in data characteristics. Using $a(i)g(i)$ as the utility metric has a theoretical grounding  [11] in the training instance sampling problem, where such a metric is shown to accelerate convergence  [6], but instead of selecting training instances, we are selecting embedding chunks to train on.
@@
Embedding pruning occurs every $T$ iterations. Empirically, $T$ is set to around one hour’s worth of training data. During pruning, pruning thresholds are computed by first sorting the current utility values and then selecting the value at the desired pruning ratio. Instead of sorting all the utility values which can take a long time, a small number of utility values are randomly sampled. After obtaining the thresholds, FIITED will calculate the number of embedding chunks crossing the threshold, i.e. previously evicted chunks with utility values larger than the threshold or previously preserved chunks with utility values now smaller than the threshold. If this number exceeds a pre-defined ratio, an enforced pruning is executed, where embedding chunks with utility below the thresholds are pruned while the rest are retained in the model.
@@
        Check if enforced pruning is needed for the pruning thresholds
        if $enforce\_pruning==True$ then
@@
Manual VS. adaptive per-chunk pruning ratios. Individual pruning ratios for each chunk can be selected either manually or adaptively according to data characteristics. For manual selection, a pruning ratio $p_{k}$, $k=0,1,...,K-1$, for each chunk is selected empirically before training. For adaptive selection, a global pruning ratio $p$ is decided by the user and indicates the average pruning ratio across all the chunks. Utility values in all chunks are sorted together to decide one global pruning threshold. Individual pruning ratios for different chunks may be bigger or smaller than $p$, and may change over time, depending on chunk utility values computed at runtime. We evaluate both approaches in the experiments (Section IV).
2D pruning. Although our current dimension optimization algorithm already captures the characteristics of embedding tables in recommender systems, there is still further potential optimization due to the uneven property of dimension importance, as discussed in Section II-B. As a result, we further harness this characteristic by introducing a dimension-level pruning method on top of the current entry-level pruning method, forming a 2D pruning system together. In dimension-level pruning, we maintain the utility values of each dimension by calculating the squared gradient sum of all embedding entries on that dimension, and prune the dimensions using a similar approach as introduced above. Specifically, we only preserve the important dimensions of all entries in the embedding table, while the unimportant ones are replaced with zero. Through this mechanism, we can further push the boundary of pruning ratios on the basis of the current design.
At the beginning of the training procedure, the entire embedding dimension is still needed to identify the hot dimensions, which we call a cold-start process. After the first dimension pruning decision is made, the embedding dimension is shrunk to the desired size, and only the preserved dimensions will be stored. Furthermore, as showcased in Section II-B, the hot dimensions, once identified, remain consistent during the whole training procedure. Hence, dimension pruning decisions only need to be made scarcely, or only during the cold-start period at the beginning of the training procedure, introducing little overhead to the current system. We offer dimension-level pruning as an optional addition to our system, where users can decide the settings based on their actual needs and the characteristics of the embedding table used during deployment.
III-C System design
Although fine-grained in-training dimension pruning has multiple benefits, it faces one practical challenge to realize memory saving during training: a straightforward implementation will result in many small fragments of free memory. To this end, we propose a new Virtually Hashed Physically Indexed (VHPI) embedding table design adapted from AdaEmbed [11], as illustrated in Figure 6. The original system design in AdaEmbed becomes a special case where the number of chunks is 1.
System Components. The design mainly consists of a hash table, an embedding table, and a chunk address manager. (1) VHPI hash table. Each hash table entry contains addresses of $K$ embedding chunks, $K$ utility values, and a bit mask that indicates whether the chunks have been pruned. Since $K$ is usually chosen to be small, the size of an entry is also small, which allows the hash table to have a large number of entries and a low collision rate. (2) VHPI embedding table. The embedding table stores embedding vector chunks coming from all sparse features. The size of the embedding table is set to a desired amount decided by the user, i.e., the pruning rate multiplied by the size of embeddings in an unpruned model. (3) Chunk address manager. It manages free addresses in the embedding table and maintains a free address stack which is updated with newly available chunk locations whenever a chunk is evicted during pruning.
Operations. (1) Embedding access. To fetch an embedding, the sparse feature ID and the feature value are together hashed to obtain the index of an entry in the VHPI hash table. Unpruned chunks are identified by checking the $K$-bit mask, and their addresses are obtained from the entry. Embedding chunks are then fetched from the embedding table according to the addresses. (2) Embedding eviction. When an embedding chunk needs to be evicted, its address is passed to the chunk address manager, who pushes the address into the free address stack, and the $K$-bit mask in the hash table is updated accordingly. (3) Embedding allocation. To allocate an embedding chunk, the chunk address manager fetches the next available chunk location, which is then stored in the hash table entry. Embeddings at the allocated location are initialized and the $K$-bit mask in the hash table is updated.
@@ III-C System design
Figure 7: Per-chunk pruning ratio settings for industry models. Figure legends show the average pruning ratio for each setting. The number of chunks varies from 4 to 8.
III-D Overhead optimization
A major challenge in the system design described in Section III-C is that a naïve implementation approach will cause significant performance overhead. This overhead mainly comes from the following sources: (1) Hash table look-up. In FIITED, we need an extra mapping process to convert input indices to their actual addresses in the embedding table, or an all-zero dummy embedding if the embedding is currently pruned. (2) Multi-chunk embedding table look-up. After chunk addresses of the sparse features have been retrieved from the hash table, embedding vectors of different chunks need to be fetched from corresponding embedding tables and then aggregated to form the entire embedding vector. (3) Utility value update. After each backward process during training, the utility value of each accessed entry needs to be updated. This involves the coalescing of gradients and calculation of utility values specified in Equation  1. (4) Pruning overhead. In a pruning process, first, a pruning threshold has to be computed by sorting the sampled utility values and then selecting the value at a desired pruning ratio, as described in  III-B. Although this overhead is already reduced by computing the threshold only on a sampled set of the whole embedding table, this process can still be expensive. Moreover, as pruning thresholds may vary between different chunks, the sampling-sorting process needs to be done for each individual chunk, further aggravating the overall time cost. After a threshold is selected, modifications need to be made on the VHPI hash tables for each chunk, evicting entries falling below the threshold and bringing back newly emerging entries.
In order to tackle the issues addressed above, we propose a set of overhead reduction methods using pipeline overlapping and parallelism techniques to optimize the performance of our system. In particular, we first use parallelism to reduce the overhead caused by the introduction of multiple chunks in our system, including embedding table look-up and pruning process. Then we further reduce the overhead by overlapping the utility value update stage of the previous batch with the forward process of the next batch. Since the calculation of utility value only requires previous frequency and gradient information, it can be overlapped with hash table look-up and embedding table look-up, reducing overhead from both sources. Furthermore, a pre-fetching technique can also be implemented in our system, where hash table look-up and embedding-table look-up can be wrapped in a dataloader to utilize the pre-fetching scheme, in which case the overhead can be further reduced by pre-fetching embeddings from in the next batch during the current training iteration.
TABLE I: Models and datasets used for evaluation.
Model Dataset Embedding Table Size
DLRM Criteo Kaggle 4.02GB
DLRM Criteo Terabyte 11.71GB
DLRM Avazu 1.13GB
DLRM MovieLens-20M 0.03GB
Model A (private) private 180GB
Model B (private) private 430GB
TABLE II: Maximum embedding table reduction on public datasets.
(ESAPN is only applicable to the MovieLens-20M dataset.)
Dataset AdaEmbed ESAPN FIITED (Manual) FIITED (Dynamic)
Criteo Kaggle 50×50$\times$50 × - $\boldsymbol{100\times}$ (2×2$\times$2 × improvement) $\boldsymbol{100\times}$ (2×2$\times$2 × improvement)
Criteo Terabyte 66.7×66.7$\times$66.7 × - $\boldsymbol{100\times}$ (1.5×1.5$\times$1.5 × improvement) $\boldsymbol{100\times}$ (1.5×1.5$\times$1.5 × improvement)
Avazu 100×100$\times$100 × - 200×200$\times$200 × $\boldsymbol{400\times}$ (4×4$\times$4 × improvement)
MovieLens-20M 1.8×1.8$\times$1.8 × 2.1×2.1$\times$2.1 × $\boldsymbol{2.7\times}$ (1.3×1.3$\times$1.3 × improvement) $\times2$
@@ III-D Overhead optimization
Figure 8: Validation accuracy under different pruning ratios (62.5%, 75%, 87.5%, 93.75%, 99%) on the Criteo Kaggle dataset. Mixed Dimension by design cannot reach 99% pruning ratio.
Figure 9: Model quality vs. pruning ratio on two industry models.
IV Experimental Results
IV-A Methodology
Baselines. We compare FIITED with the following baselines: (1) AdaEmbed [11], a state-of-the-art in-training embedding pruning method that prunes entire embedding rows during training, as described in Section III-A; (2) ESAPN [14], a state-of-the-art in-training EDS method that selects embedding dimensions by training additional policy networks that decide to enlarge or shorten embedding dimensions during training; (3) Mixed Dimension (MD) embeddings [5], a state-of-the-art pre-training EDS method that selects per-feature embedding dimensions prior to training based on access frequency; (4) last but not least, the original unpruned DLRM model with uniform embedding dimensions, which will be denoted simply as DLRM in the comparison. Baseline selection criteria: Since the main contribution of FIITED is to save memory during training, we selected 2 in-training embedding size optimization methods that can achieve a similar effect, in addition to 1 pre-training method that reduces the model size before training. We did not select post-training EDS methods  [1, 32, 10] or other in-training EDS methods  [26, 33, 9, 2], because they cannot reduce training memory usage, as discussed in Section V.
Models and datasets. We evaluate FIITED on public models and datasets as well as private industry ones. A summary of the models and datasets can be found in Table I. (1) Public models and datasets: We adopt an open-source DLRM framework 111 https://github.com/facebookresearch/dlrm to train a DLRM [17] on Criteo Kaggle Display Advertising Challenge Dataset 222 https://ailab.criteo.com/ressources, Avazu Click-Through Rate Prediction Dataset 333 https://www.kaggle.com/c/avazu-ctr-prediction and Criteo Terabyte Dataset  [^4] https://labs.criteo.com/2013/12/download-terabyte-click-logs. We also modified this framework to train another open-source DLRM 555 https://github.com/zgahhblhc/ESAPN used by ESAPN [14] on the MovieLens-20M dataset 666 https://grouplens.org/datasets/movielens/20m. On the MovieLens-20M dataset, following ESAPN, we converted the multi-classification problem to a binary classification problem by viewing 4-star and 5-star reviews as positive, and others as negative. (2) Industry models and datasets: Two production models are used in the evaluation. For ease of experimentation, both models are shrunk to 1/4 size by reducing the hash size of all sparse features by 75%. The reduced models, named Model A and Model B, have around 180GB and 430GB sparse feature embeddings, respectively, and both contain hundreds of sparse features. They are evaluated on data generated by real-world applications. Ten consecutive days of data are used as the training set, and the following one day’s data are used as the evaluation set.
Implementation. For evaluation using the public models and datasets, we implement FIITED as well as AdaEmbed [11] on top of an open-source DLRM framework [17] based on PyTorch. Utility values are asynchronously updated with training to reduce runtime overhead, while operations for different chunks are parallelized, as described in  III-D. For evaluation on the industry models and datasets, a proof-of-concept prototype design is implemented on the company’s internal deep recommendation system code base. The prototype stores both pruned and unpruned embedding chunks, and executes pruning by setting the corresponding memory regions to zero.
Training specifications. For evaluation using the public models and datasets, the experiments are run on 1 GPU, and we compare two methods to select per-chunk pruning ratios: manually and dynamically. To manually specify the pruning ratio for each embedding chunk, a linear function is fitted to satisfy the desired average pruning ratio. We used a default sparse embedding dimension of 32 on Criteo Kaggle and Avazu datasets, 64 on Criteo Terabyte, 128 on MovieLens-20M dataset. The number of chunks $K$ is set to 2 on all datasets for FIITED.
For evaluations using the industry models and datasets, the experiments are run on 32 A100 GPUs, and the average embedding pruning ratio varies from 30% to 85%. Per chunk pruning ratios are manually chosen and shown in Figure 7. The ratios are designed to roughly resemble a power law distribution, which was assumed by previous EDS work  [5] and fits the observation that only a small number of “hot” embeddings require a long dimension  [20].$K$ varies between 4 and 8. For Model A, we also experimented with 2 different pruning ratio settings that have the same average ratio (75%), one is “steeper” with the ratios in different chunks span a larger range and the other is “flatter”, to investigate the effect of such a difference. Chunk utility is simply computed as a running average of access frequency in the experiments on industry models.
Metrics. In this work, we mainly care about (i) maximum embedding memory saving without significantly affecting model quality. This metric showcases the core ability of FIITED by reducing the embedding table size without harming its prediction quality. (ii) overhead introduced by FIITED. As a longer end-to-end training time will cause more resource consumption, it is important to ensure that the overhead of FIITED is negligible compared with the unpruned baseline, and this overhead grows at a controllable scale as the chunk number grows.
TABLE III: Maximum embedding table memory saving using FIITED-2D.
Dataset Entry Reduction Dimension Reduction Overall Reduction
Criteo Kaggle 66.7×66.7$\times$66.7 × 2×2$\times$2 × 133.3×133.$3$
Criteo Terabyte 66.7×66.7$\times$66.7 × 2×2$\times$2 × 133.3×133.$3$
Avazu 100×100$\times$100 × 8×8$\times$8 × 800×800$\times$800 ×
MovieLens-20M 1.8×1.8$\times$1.8 × 12.5×12.5$\times$12.5 × 17.8×17.8$\times$17.8 ×